HUBBLE SPACE TELESCOPE DAILY REPORT #2530 PERIOD COVERED: 0000Z (UTC) 12/30/99 - 0000Z (UTC) 01/03/00 Daily Status Report as of 003/0000Z 1.0 OBSERVATIONS SCHEDULED AND ACCOMPLISHED: 1.1 Completed Two Sets of S/C 8518 (Vehicle Disturbance Test) The Vehicle Disturbance Test {VDT} is used to characterize the HST dynamic response due to environmental disturbances during extended gyro hold intervals. Disturbances are caused by on-board mechanism motion and transient effects resulting from thermal strain. These disturbances excite the natural modes of HST structural appendages and produce vehicle dynamic response as measured by the rate gyros. Disturbance signatures are compared with signatures taken after previous servicing missions to identify changes in mechanisms or structures since the last VDT. The disturbances observed will be used as the nominal on-orbit disturbances in pointing control simulations until the next VDT is run. At least 5 orbits at +V3 sunpoint {with 90-degree solar array angle} and 3 orbits at -V1 sunpoint {with 0-degree solar array angle} are required for this test. Using two different solar array orientations allows additional identification of disturbances induced by the solar arrays. Low bandwidth attitude control gains are used to configure the HST for "nearly open-loop" dynamic response during the test. This configuration allows observation of HST disturbance response that is nearly independent of the affects of the pointing control law. No science instruments are used during this test. The test completed with no reported problems. 1.2 Completed Six Sets of STIS/CCD 8507 (CCD Darks) The Space Telescope Imaging Spectrograph (CCD) was used to obtain weekly updates to the hot pixel list. The proposal completed as planned. 1.3 Completed Five Sets of STIS/CCD 8508 (CCD Bias Monitor C8) The Space Telescope Imaging Spectrograph (CCD) was used to obtain updates to the weekly superbiases. The proposal was completed as planned. 1.4 Completed Ten Sets of STIS/CCD 8530 (SMOV3a STIS BEA Science) The Space Telescope Imaging Spectrograph (CCD) was used to obtain SMOV3a Bright Earth Avoidance time period for science exposures with the STIS CCD. Various kinds of targets are observed. During the ninth iteration of this proposal, the loss of lock described in 2.1 and HSTAR 7454 occurred. The Science Init Flag never came up. Otherwise, the proposals completed nominally. 1.5 Completed Eight Sets of FGS 8532 (White Dwarf and M-Stars in the Solar Neighborhood; Revealing Binarity) FGS-1 was used to conduct a high angular resolution survey with FGS1r to detect binarity among nearby M dwarfs and white dwarfs. During fourth, fifth and seventh iterations of this proposal, the Astrometry FGS failed the search radius limit test. HSTARs 7453, 7450, and 7457 were written. Otherwise, the remaining observations completed with no reported problems. 1.6 Completed Two Sets of STIS/CCD/MA1/MA2 8514 (Contamination Test) The Space Telescope Imaging Spectrograph (CCD/MA1/MA2) was used to investigate whether there is excessive Lyman Alpha contamination {decrease in throughput} for WFPC2 observations. The observations completed with no reported problems. 1.7 Completed Three Sets of STIS/MA1/MA2 8504 (SM3A MAMA Turn-on) The Space Telescope Imaging Spectrograph (MA1/MA2) was used to turn on MAMA high voltages and take darks three times per day. The observations completed nominally. 1.8 Completed WF/PC-22 8498 (Internal Monitor - Post-TEC Turn On) The WF/PC-2 was used to obtain routine internal observations in order to monitor basic instrument health, and optical alignment (K-spots, INTFLATS, VISFLATS, UVFLATS, bias, Darks). The observations completed nominally. 1.9 Completed Two Sets of WF/PC-2 8491 (Transition to hold, Decon and Cool Down Procedure) This program covers the transition from Protect Safe Mode to Hold Mode, then performs a Protect Decontamination. This is followed by a series of UV monitors and focus checks at -43C, then a special decon to take the instrument to -88C. The program completed nominally. 1.10 Completed STIS/MA1 8510 (FUV MAMA Darks) The Space Telescope Imaging Spectrograph (MA1) was used to determine the health of the FUV MAMA detector by monitoring the background count rate. The observations completed with no reported problems. 1.11 Completed Four Sets of STIS/MA2 8509 (NUV MAMA Darks) The Space Telescope Imaging Spectrograph (MA2) was used to determine the health of the FUV MAMA detector by monitoring the background count rate. The observations completed with no reported problems. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions: 47 Successful: 47 Scheduled Re-acquisitions: 16 Successful: 16 HSTARs 7450, 7453 and 7457 document cases where a successful acquisition was followed by the astrometry FGS-1 failing the search radius limit test. The proposal detailed in was affected. Per HSTAR 7454, there was a loss of lock at 002/005623Z. The proposal described in 1.4 was affected. 2.2 FHST Updates: Scheduled: 137 Successful: 137 Although the roll delay update at 365/032635Z, using FHST-2, succeeded, the target suppression B-2 flag was out-of-limits (usually indicating failure). This occurred with the FHST assigned to both electronics A and B. The mnemonic remained out-of-limits until the tracker was commanded closed for the type 2 slew. HSTAR 7451 was written. 2.3 Operations Notes: Servicing Mission Orbital Verification (SMOV) continues. The correctable SSR EDAC error counter was reset six times per ROP SR-1A. During the weekend, the SSR science playback pointer was set four times as directed by ROP SR-3. Per NS-5, the SI C&DH error counter was reset (for Exec 101, ENGREP 0) twice. Four TTRs were written for transmitter scheduling errors during the long weekend period: for day 364 there was seven errors; for day 365, ten errors: for day 001, 19 errors: and for day 002, 24 errors. Another TTR was written for a 6 minute, unrecoverable loss of 32k engineering data due to an HGA spline and an MA transmitter switch during the support at 365/0304Z.. At 365/105725, the coarse sun sensor #1 temperature (GCSS1T) flagged out-of-limits low with a value of -45.1 deg. C for one six-second sample. HSTAR 7452 was written. The Y2K rollover occurred with no problems reported. There was a STIS MCE-1 reset at 001/172502Z while the low voltage was on and while outside any SAA interval. The STIS flight software error counter was reset at 001/1832Z per ROP-NS-12. The STIS MAMA-1 was recovered at 003/0959Z, the time that the high voltage was commanded on via the SMS. The FHST-21 temperature flagged out-of-limits high at 002/110225Z for one telemetry sample. This was documented in HSTAR 7455. At 002/172859Z, the FHST-21 shelf temperature flagged out-of-limits high (12.0 deg. C) several times during an eight minute time period. HSTAR 7456 was written. The limit was raised to 13.0 deg. C per an operations note. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Continuation of SMOV and normal science observations and calibrations. /CAW